OpenStudy (anonymous):

Please teach me the steps to solve this problem Suppose sin Alpha= 4/5, 0

OpenStudy (freckles):

do you remember the double angle identity for sine? \[\sin(2 \alpha)=2 \sin(\alpha)\cos(\alpha)\] also draw a right triangle in the appropriate quadrant. 

OpenStudy (freckles):

find the adjacent side

OpenStudy (freckles):

you want to find the adjacent side so you find cos(alpha)

OpenStudy (anonymous):

25-16=sqr9=3 so 2(4/5)(3/5)?